kimberly bought a bag at dicount of 20%. if the original price of the bag was $150, how much did kimberly pay?
Answer by garettbuff(18)

150=100%
100%-20%=80%
150
×.8
---
120=x


Or 1/5 =20% so 150/5=30
30=20%
100%-20%=80%
150-30=120
80%=120